OpenStudy (batman31):

how would I convert 18/sqrt3 and 10/sqrt3 to simplest radical form?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[\frac{18}{\sqrt3}=\frac{18}{\sqrt3}\times\frac{\sqrt3}{\sqrt3}=\frac{18\sqrt3}{3}=6\sqrt3\] second one is similar

OpenStudy (batman31):

I get how you did the first one now, but I dont quite understand how to apply this to the second one yet.

OpenStudy (batman31):

10 isn't divisible by three, and I think thats throwing me off @misty1212

OpenStudy (batman31):

does that mean it's already in simplest radical form?

OpenStudy (anonymous):

You can just leave out the final step if it doesn't divide evenly

OpenStudy (batman31):

so it already is in simplest radical form?

OpenStudy (anonymous):

\( \frac{18}{\sqrt3}=\frac{18}{\sqrt3}\times\frac{\sqrt3}{\sqrt3}=\frac{18\sqrt3}{3}= \color{red}{6\sqrt3} \) If you can't divide, leave out only the red part.

OpenStudy (batman31):

so (10sqrt3)/3 is simplest radical form?

OpenStudy (anonymous):

That's pretty much the agreed 'most simple' form.

OpenStudy (anonymous):

Even if it looks less simple, it's the most commonly accepted way to simplify it.

OpenStudy (batman31):

thanks, can you medal misty for me since you both helped?

OpenStudy (anonymous):

sure, no problem